Thursday, Mar. 03: SWING Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans


  • Complete energy measurement with series of Hall-A magnet (mapping)
- HALOG [1]
  • RF recovery done, Prepare the BCM calibration (not happened due to delay restoring the beam)
  • Set up beam through Compton chicane

Specific mission

  • Note:

1/ Set up beam through Compton chicane

- Ciprian, Greg and Yves  are tuning the beam through Compton
- HALOG [2] - Before tune the beam, observe the beam current (how much actually we can get tonight but only up to 30uA) - Spoke to Ciprian/Greg/Dave and they gave a specific monitoring instruction (shift crew will post everything). - We will give a try ...

2/ Expect some production runs through Compton

3/ Agreement with Hall-D and Talked PD to set up the BCM calibration tomorrow

- Crucial: high beam current 30uA with FC & empty target 

Instruction: DOC1DOC2
- Experts on call (Dave/Julie/Thir/...) - Calo. HV off

BPM/Spot++ Check

  • After beam lost for considerable time (like beam study, pass change,...)
  • Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++